OpenStudy (anonymous):

What is the slope and y-intercept of the graph of 2x + 3y = 18? slope = -2; y-intercept = (0, 18) slope = negative 2 over 3; y-intercept = (0, 6) slope = negative 3 over 2; y-intercept = (0, 2) slope = 6; y-intercept = (0, negative 2 over 3)

OpenStudy (eyust707):

Still need help?

OpenStudy (anonymous):

Yes, Please.

OpenStudy (anonymous):

dividing by 3 on both sides we have (2/3)x+y=6 This implies y=(-2/3)x+6=mx+c where m=-2/3 and c=6. What can you say then?.@kaylamarie081

OpenStudy (anonymous):

@kaylamarie081

OpenStudy (ankitshaw):

make the eqn of the line in the form of y= mx+c ,where m is the slope and c is the y intercept.

OpenStudy (anonymous):

Is my answer choice B?

OpenStudy (anonymous):

@kaylamarie081. correct